OpenStudy (anonymous):

P = 2(l + w) for l. What are the missing values in the table? P w l 14 2 ? 22 8 ?

OpenStudy (anonymous):

p=2l+2w ; 2l=p-2w ; l=p/2-w

OpenStudy (anonymous):

\[l = (P/2)-w\] now substitute the values of 'P' and 'w'